OpenStudy (anonymous):

-tan ^ 2 x+ sec ^2 x = 1

OpenStudy (agent0smith):

First change everything into sine and cosine\[\Large -\frac{ \sin^2 x }{ \cos^2 x }+ \frac{ 1 }{ \cos ^2 x }\]now you can add the fractions since they have like denominators \[\Large \frac{- \sin^2 x +1 }{ \cos^2 x }\]maybe change the order \[\Large \frac{1- \sin^2 x }{ \cos^2 x }\]Does that top part look familiar?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

1 - sin^2 (x) = cos^2 (x) so then it would be cos^2 (x) / cos^2(x)

OpenStudy (anonymous):

Right

OpenStudy (agent0smith):

And cos^2 x/ cos^2 x is equal to...

OpenStudy (anonymous):

sine ?

OpenStudy (agent0smith):

No... a number divided by itself equals 1. What then must this be equal to? \[\Large \frac{ \cos^2 x }{ \cos^2 x }= ???\]